# NaviAgent: Bilevel Planning on Tool Dependency Graphs for Function Calling

### v1

@@ 3.2.4 Graph Evolution
The TDHG in real-world environments is inherently dynamic, evolving as new tools are introduced, obsolete ones are deprecated, and usage patterns shift. As shown at the bottom of Figure 1: At time $T-1$, selectable paths are relatively uniform (similar line shades), indicating multiple equally viable routes. By $T$, the Navigator has learned to prefer a more optimal path (the top, darker line). At $T+1$, when the upper-left API becomes unavailable, the Navigator adapts by recomputing and selecting the lower route as the new optimal path. This demonstrates the Navigator’s ability to learn from feedback and flexibly adjust its planning in response to runtime changes. To systematically support such adaptability, we design a graph evolution framework with three key mechanisms:
@@ 3.3 Dynamic Execution & Path Recombination
Robust and adaptive toolchain orchestration is achieved through a bilevel dynamic planning framework, in which the Decider manages action selection and the Navigator is responsible for toolchain planning.
Collaboration Workflow.
The case in Figure 2 demonstrates the workflow: given a multi-intent query, the Decider first decomposes it into two sub-tasks, one that can be answered directly and another that requires tool retrieval. For the sub-task requiring tool use, the Decider employs a BGE-based retriever to identify three target APIs (highlighted as red solid nodes), and invokes the Navigator to perform backward graph search, yielding three candidate subgraphs $\mathcal{G}^{\prime}$ (composed of red edges). These are then returned to the Decider, which executes the corresponding tool call. More detailed cases can be found in the Appendix C.
Figure 2: Case Study of the Collaboration Workflow.
@@
If a tool call action fails, for example, due to an unavailable or malfunctioning API, the Decider will not terminate or rigidly adhere to the original plan. Instead, it selects the tool retrieval action, triggering a new round of collaboration with the Navigator. The Navigator then performs graph search on the current TDHG to identify an alternative toolchain. The Decider subsequently executes the new plan, and this process can be repeated as necessary until the task is completed or determined to be infeasible. This dynamic path recombination allows the agent to adaptively recover from failures and maximize task completion in complex tool environments.
@@ 4.1 Experimental Settings
We compare our approach with several representative baselines, including ReAct yao2023react (alternates reasoning steps and tool calls), ToolLLM qin2023toolllm (DFSDT-based tool planning and execution), and $\alpha$-UMI shen2024small (multi-agent framework with modular tool-use stages handled by lightweight LLMs). Experiments are conducted across multiple foundation models, including open-source models (Qwen2.5-14B yang2024qwen2 , Qwen2.5-32B tahmid2024qwen2 , DeepSeek-R1-Distill-Qwen-32B(DeepSeek-R1-32B) guo2025deepseek ) and closed-source models (DeepSeek-V3 liu2024deepseek , GPT-4o hurst2024gpt ), as well as a fine-tuned lightweight model (Qwen2.5-14B).
@@ 4.2 Results
Model Method Easy Medium Hard All TCR TSR Steps TCR TSR Steps TCR TSR Steps TCR TSR Steps Qwen2.5-14B ReAct 32.4 26.3 3.52 24.5 16.8 3.67 24.8 20.0 3.64 27.1 20.6 3.61 ToolLLM 56.1 30.4 4.01 53.8 19.7 4.02 38.1 11.4 4.06 51.0 21.3 4.03 $\alpha$-UMI 77.7 39.2 5.53 77.9 25.0 5.88 67.6 13.3 6.07 75.5 26.9 5.74 NaviAgent 64.2 50.7 4.32 60.6 32.7 4.32 61.9 23.8 4.47 62.0 36.4 4.34 Qwen2.5-32B ReAct 33.1 25.0 3.50 35.6 24.5 3.60 30.5 19.0 3.95 33.6 23.4 3.63 ToolLLM 40.5 31.8 3.67 48.6 30.3 3.85 49.5 23.8 4.10 46.2 29.3 3.83 $\alpha$-UMI 78.4 49.3 5.66 78.8 26.0 6.02 77.1 22.9 6.58 78.3 32.8 5.94 NaviAgent 88.5 60.8 4.33 82.2 42.8 4.70 80.0 26.7 5.11 83.7 44.9 4.60 Deepseek-V3 ReAct 46.6 36.5 3.52 58.7 38.5 3.50 48.6 23.8 3.74 52.5 34.5 3.54 ToolLLM 56.2 47.4 3.80 58.8 30.0 3.92 29.7 24.8 3.90 51.3 34.4 3.86 $\alpha$-UMI 80.8 59.7 5.95 89.4 32.9 5.95 73.0 29.5 6.64 82.9 40.7 6.06 NaviAgent 99.3 72.3 4.40 96.6 49.0 4.46 96.8 44.8 5.10 98.3 55.5 4.55
@@
As shown in Table 1 and Figure 3, NaviAgent consistently achieves the highest TSR across all foundation models and task complexities, with absolute values of 36.4% on Qwen2.5-14B, 44.9% on Qwen2.5-32B, and 55.5% on Deepseek-V3. Compared to the average performance of the baselines on all tasks, NaviAgent achieves substantial gains of 13.47, 16.4, and 19.0 percentage points on Qwen2.5-14B, Qwen2.5-32B, and Deepseek-V3, respectively. Meanwhile, its execution steps remain close to those of the most efficient baseline, with differences typically within one step, thereby maintaining a strong balance between solution quality and execution efficiency.
@@
NaviAgent achieves an average TSR improvement of over 10 percentage points compared to $\alpha$-UMI, tthe strongest among the three baselines, across all difficulty levels, with the most significant gain of 15.33 percentage points on Deepseek-V3 for Hard tasks. We further observe that the relative drop in TSR from Easy to Hard tasks is substantially smaller for NaviAgent than for most baselines, particularly on larger foundation models. For example, on Deepseek-V3, NaviAgent’s TSR decreases by only 38.0% from Easy to Hard, while ToolLLM and $\alpha$-UMI experience drops of 47.7% and 50.6%, respectively.
@@
Notably, with supervised fine-tuning, the smaller Qwen2.5-14B model achieves performance comparable to the larger 32B model (TCR 81.8% vs 83.7%, TSR 49.5% vs 44.9%, see Figure 4 and Table 2, D+N(Heur) row), indicating that fine-tuning can effectively close the gap between model sizes.
@@ 4.3 Ablation Study
Multi-Path Decider Analysis.
@@
We analyze the execution paths taken by only Decider in successful cases on ToolBench, focusing on the Deepseek-V3 model. Specifically, we categorize the proportion of cases resolved via Clarification (intent clarification to seek additional details from the user), Re-retrieval (recovering from initial tool retrieval failures by invoking alternative APIs), and Normal (tasks completed successfully in a single attempt without clarification or re-retrieval). Results are summarized in Figure 5, demonstrating that multi-path planning enables robust error recovery and flexible intent handling, contributing to overall performance gains.
Figure 5: Comparison of TSR Distribution Between Multi-Path Decider and Baselines.
Graph-Encoded Navigator Analysis.
Table 2 indicates that introducing the Navigator module consistently improves TSR across all tasks and model sizes, with our heuristic search strategy providing additional gains, achieving an average TSR improvement of 2.4 points over Alpha-Beta search. Notably, these benefits become more pronounced as model capacity increases. On hard tasks, the addition of the Navigator yields relative TSR improvements of 27.3% for Deepseek-V3 and 36.2% for GPT-4o. This demonstrates that graph-based planning and efficient search more effectively unlock the reasoning and compositional capabilities of larger models, especially for complex toolchain tasks. Consistent results are also observed on API-Bank (see Table 5). Additional statistics on TDHG structure and link prediction are provided in Table 6.
@@
Model Method Easy Medium Hard All TCR TSR Steps TCR TSR Steps TCR TSR Steps TCR TSR Steps Qwen2.5-14B Decider 54.7 41.2 4.35 58.2 29.3 4.30 66.7 21.0 4.55 59.0 31.2 4.36 D+N(Alpha) 58.8 48.0 4.21 61.5 31.7 4.33 53.3 16.2 4.28 58.8 33.4 4.27 D+N(Heur) 64.2 50.7 4.32 60.6 32.7 4.32 61.9 23.8 4.47 62.0 36.4 4.34 Qwen2.5-32B Decider 84.5 52.0 4.32 81.3 40.9 4.83 90.5 24.8 5.70 84.4 40.8 4.74 D+N(Alpha) 83.1 51.4 4.31 85.1 41.3 4.87 80.0 31.4 5.04 83.3 42.3 4.68 D+N(Heur) 88.5 60.8 4.33 82.2 42.8 4.70 80.0 26.7 5.11 83.7 44.9 4.60 Deepseek-R1-32B Decider 94.6 40.5 5.01 92.8 34.1 5.49 93.3 27.6 5.29 93.5 34.7 5.27 D+N(Alpha) 92.6 51.4 4.96 93.3 38.0 5.17 91.4 21.9 5.60 92.6 38.6 5.13 D+N(Heur) 93.2 51.4 4.98 92.3 38.0 5.33 88.6 33.3 5.27 91.8 41.2 5.18 Deepseek-V3 Decider 99.9 71.6 4.26 99.0 47.1 4.88 99.9 35.2 4.95 99.6 52.3 4.62 D+N(Alpha) 93.2 71.6 4.26 95.7 50.5 4.52 93.3 33.3 4.64 94.4 53.4 4.42 D+N(Heur) 99.3 72.3 4.40 96.6 49.0 4.46 96.8 44.8 5.10 98.3 55.5 4.55 GPT-4o Decider 99.8 67.6 4.05 98.6 43.9 4.60 99.0 35.9 4.95 99.1 49.7 4.42 D+N(Alpha) 99.9 76.4 4.08 99.5 45.3 5.24 98.1 41.4 5.59 99.3 54.4 4.78 D+N(Heur) 99.9 72.1 4.05 95.7 50.0 4.78 99.0 48.9 4.96 97.8 56.8 4.52 Qwen2.5-14B(SFT) Decider 77.7 55.4 4.80 78.8 46.2 4.78 75.2 30.5 5.57 77.7 45.6 4.91 D+N(Alpha) 85.8 64.9 4.48 78.4 39.9 4.59 78.1 39.0 5.26 80.7 47.7 4.67 D+N(Heur) 83.8 64.2 4.68 81.7 49.0 4.73 79.0 29.5 5.66 81.8 49.5 4.84
Table 2: Impact of Navigator and Search Algorithms on ToolBench. Decider refers to using only the decider module. D+N(Alpha) denotes Decider + Navigator with Alpha-Beta search. D+N (Heur) denotes Decider + Navigator with our proposed heuristic search. Metrics are reported as in Table 1.
@@ 5 Conclusion
We present NaviAgent, a bilevel planning framework for compositional tool reasoning that enhances large language models’ planning and execution abilities. By combining a decider and a graph-based navigator with heuristic search, NaviAgent delivers robust improvements in task completion and success rates across diverse tasks and model sizes, while maintaining competitive efficiency. Experiments on ToolBench and API-Bank demonstrate that NaviAgent is particularly effective for complex, multi-tool tasks and larger models, consistently achieving strong TSR gains.
Limitations.
Despite its strengths, NaviAgent has several limitations. First, integration challenges from heterogeneous tool interfaces remain to be improved, which could be mitigated by adopting unified protocols like MCP. Second, the single-agent design limits collaboration and generalization in broader or multimodal tasks. Finally, although NaviAgent incorporates adaptive graph evolution, highly dynamic or noisy environments could still challenge robust graph construction and heuristic search. Addressing these issues will be important for future work.
@@ B.1 HGT Network
B.2 Training Objective
This section provides the full definitions of the loss functions and training strategies used for graph optimization.
Cross-Entropy Loss.
The cross-entropy loss encourages the model to match predicted link probabilities with observed edge weights:
Adaptive Margin Loss.
To further enhance representation discrimination, we introduce an adaptive margin loss that dynamically assigns larger separation intervals to node pairs with higher edge weights (i.e., $\tilde{w}_{uv}\rightarrow 1$), encouraging the model to focus on learning critical dependencies. For each positive edge $(u,v)^{+}\in\mathcal{E}^{+}$, $k$ negative edges $\{(u_{j},v)\}_{j=1}^{k}$ are sampled to construct positive and negative pairs for the margin loss.
$m_{uv}=m_{0}\left(1+\sigma(\tilde{w}_{uv})\right)$ (17)
where $m_{0}$ is a base margin, $\sigma(\cdot)$ denotes the sigmoid function, $\tilde{w}_{uv}$ is the statistical edge weight, $s(u,v)$ measures the embedding similarity, $(u,v)^{+}$ represents a positive edge, $(u_{j},v)^{-}$ denotes a negative sample, and $[\cdot]_{+}$ is the hinge function, and $\mathcal{E}^{+}$ is the set of positive edges.
Hybrid Loss.
The final training objective is a weighted sum of the two losses:
$\mu_{t}=\mu_{0}\cdot\gamma^{t},\ \gamma\in(0,1)$ (19)
where $\mu_{t}$ is the weight for the cross-entropy loss at epoch $t$, $\mu_{0}$ is the initial weight, and $\gamma$ is a decay factor controlling the rate at which the contribution of the cross-entropy loss decreases over training.
B.3 Graph Search Algorithm

### v2

@@ 3.2.4 Graph Evolution
The tool world is inherently dynamic, evolving as new tools are introduced, obsolete ones are deprecated, and usage patterns shift. As shown at the right of Figure 2: At time T1T-1, selectable paths are relatively uniform (similar line shades), indicating multiple equally viable routes. By $T$, the TWNM has learned to prefer a more optimal path (the top, darker line). At $T+1$, when the upper-left API becomes unavailable, the TWNM adapts by recomputing and selecting the lower route as the new optimal path. This demonstrates the TWNM’s ability to learn from feedback and flexibly adjust its planning in response to runtime changes. To systematically support such adaptability, we design a graph evolution framework with three key mechanisms:
@@ 3.3 Dynamic Execution & Path Recombination
Robust and adaptive toolchain orchestration is achieved through a bilevel dynamic planning framework, in which the agent manages action selection and the TWNM is responsible for toolchain planning.
NaviAgent Workflow.
When a user query arrives, NaviAgent decides whether it can respond directly, clarify the user’s intent, or rely on external tools. For more complex queries, NaviAgent decomposes the task into sub‑tasks and categorizes them into two types: those that can be answered or clarified immediately, and those that demand toolchain retrieval. Unlike traditional agents that fetch tools sequentially, NaviAgent searches the existing tool dependency graph for a task‑relevant subgraph and selects a feasible execution path for subsequent execution. More detailed cases can be found in the Appendix C.
@@
During execution, if a tool execution action fails due to an an API is unavailable or malfunctioning, the agent switches from execution to toolchain retrieval and invokes its TWNM module. TWNM searches the current tool dependency graph to recombine nodes and identify an alternative toolchain, which the agent then executes. This adaptive loop can be repeated until completion or infeasibility, enabling dynamic path recombination that improves robustness and task success in complex tool environments.
@@ 4.1 Experimental Settings
The evaluation considers frameworks for real-world tool invocation, where managing large tool sets and enabling autonomous planning are critical. We select representative baselines in three major categories: (i) ReAct-based single-agent frameworks, where ReAct yao2023react serves as the foundational approach alternating reasoning and tool use; (ii) enhanced single-agent frameworks, where ToolLLM qin2023toolllm incorporates DFSDT-based planning with a dynamic backtracking mechanism; and (iii) multi-agent frameworks, where $\alpha$-UMI shen2024small organizes modular tool-use stages via lightweight LLMs. Experiments are conducted across multiple foundation models, including open-source models (Qwen2.5-14B yang2024qwen2 , Qwen2.5-32B tahmid2024qwen2 , DeepSeek-R1-Distill-Qwen-32B(DeepSeek-R1-32B) guo2025deepseek ) and closed-source models (DeepSeek-V3 liu2024deepseek , GPT-4o hurst2024gpt ), as well as a fine-tuned lightweight model (Qwen2.5-14B).
@@ 4.2 Results
Model Method Easy Medium Hard All TCR TSR Steps TCR TSR Steps TCR TSR Steps TCR TSR Steps Qwen2.5-14B ReAct 32.4 26.3 3.52 24.5 16.8 3.67 24.8 20.0 3.64 27.1 20.6 3.61 ToolLLM 56.1 30.4 4.01 53.8 19.7 4.02 38.1 11.4 4.06 51.0 21.3 4.03 $\alpha$-UMI 77.7 39.2 5.53 77.9 25.0 5.88 67.6 13.3 6.07 75.5 26.9 5.74 NaviAgent 64.2 50.3 4.18 60.1 32.3 4.38 61.1 22.4 4.68 61.6 35.8 4.38 Qwen2.5-32B ReAct 33.1 25.0 3.50 35.6 24.5 3.60 30.5 19.0 3.95 33.6 23.4 3.63 ToolLLM 40.5 31.8 3.67 48.6 30.3 3.85 49.5 23.8 4.10 46.2 29.3 3.83 $\alpha$-UMI 78.4 49.3 5.66 78.8 26.0 6.02 77.1 22.9 6.58 78.3 32.8 5.94 NaviAgent 88.1 61.1 4.29 81.7 41.5 4.60 79.4 30.8 5.31 83.2 45.4 4.66 Deepseek-V3 ReAct 46.6 36.5 3.52 58.7 38.5 3.50 48.6 23.8 3.74 52.5 34.5 3.54 ToolLLM 56.2 47.4 3.80 58.8 30.0 3.92 29.7 24.8 3.90 51.3 34.4 3.86 $\alpha$-UMI 80.8 59.7 5.95 89.4 32.9 5.95 73.0 29.5 6.64 82.9 40.7 6.06 NaviAgent 97.9 71.8 4.40 96.3 48.5 4.45 97.0 44.9 5.19 97.0 55.2 4.60
@@
As shown in Table 1 and Figure 3, NaviAgent consistently achieves the highest TSR across all foundation models and task complexities, with absolute values of 35.8% on Qwen2.5-14B, 45.4% on Qwen2.5-32B, and 55.2% on Deepseek-V3. Compared to the average performance of the baselines on all tasks, NaviAgent achieves substantial gains of 12.9, 16.9, and 18.7 percentage points on Qwen2.5-14B, Qwen2.5-32B, and Deepseek-V3, respectively. Meanwhile, its execution steps remain close to those of the most efficient baseline, with differences typically within one step, thereby maintaining a strong balance between solution quality and execution efficiency. Consistent performance is also observed in our real-world API tests, with detailed results provided in Appendix F.1.
@@
NaviAgent achieves an average TSR improvement of over 10 percentage points compared to $\alpha$-UMI, tthe strongest among the three baselines, across all difficulty levels, with the most significant gain of 15.4 percentage points on Deepseek-V3 for Hard tasks. We further observe that the relative drop in TSR from Easy to Hard tasks is substantially smaller for NaviAgent than for most baselines, particularly on larger foundation models. For example, on Deepseek-V3, NaviAgent’s TSR decreases by only 37.5% from Easy to Hard, while ToolLLM and $\alpha$-UMI experience drops of 47.7% and 50.6%, respectively.
@@
Notably, with supervised fine-tuning, the smaller Qwen2.5-14B model achieves performance comparable to the larger 32B model (TCR 81.2% vs 83.2%, TSR 51.3% vs 45.4%, see Figure 4 and Table 2, D+N(Heur) row), indicating that fine-tuning can effectively close the gap between model sizes.
@@ 4.3 Ablation Study
Effect of base Components.
@@
We analyze the NaviAgent (Base) configuration, focusing on its four‑dimensional decision space in successful ToolBench cases with the Deepseek‑V3 model. Specifically, we categorize the proportion of cases resolved via Clarification (intent clarification to seek additional details from the user), Re-retrieval (recovering from initial toolchain retrieval failures by invoking alternative APIs), and Normal (tasks completed successfully in a single attempt without clarification or re-retrieval). Results are summarized in Figure 5, demonstrating that the four‑dimensional decision space of the agent enables robust error recovery and flexible intent handling, contributing to overall performance gains.
Figure 5: Comparison of TSR Distribution Between NaviAgent(base) and Baselines.
Effect of TWNM Components.
Table 2 shows clear gains from each design choice. Compared with the Base (agent only), NaviAgent (DynamicH) improves TSR by +11.8 points on average, confirming the value of graph‑based planning with search. Dynamic graphs further outperform static ones on hard tasks (e.g., +5.1 on Qwen2.5‑32B, +2.0 on GPT‑4o), and replacing Alpha‑Beta with heuristic search brings the best results, adding 2–3 points on all tasks and about 8 points on hard cases for large models such as Deepseek‑V3 and GPT‑4o, highlighting that dynamic graph planning and efficient heuristic search are crucial for unlocking the reasoning and compositional potential of frontier models. Consistent results are also observed on API-Bank (see Table 7). Additional statistics on tool graph structure and link prediction are provided in Table 8.
@@
Model Method Easy Medium Hard All TCR TSR Steps TCR TSR Steps TCR TSR Steps TCR TSR Steps Qwen2.5-14B Base 46.4 36.0 5.38 50.5 22.9 5.39 62.0 16.3 5.76 51.8 25.6 5.47 Static+A 57.3 43.7 4.37 61.2 29.0 4.54 53.0 14.0 4.59 58.1 30.3 4.50 Dynamic+A 58.8 48.0 4.31 61.5 31.7 4.49 53.3 16.2 4.61 58.8 33.4 4.46 Dynamic+H 64.2 50.3 4.18 60.1 32.3 4.38 61.1 22.4 4.68 61.6 35.8 4.38 Qwen2.5-32B Base 77.7 47.7 5.42 75.8 32.7 6.00 86.9 19.0 7.04 78.9 34.4 6.05 Static+A 82.8 50.7 4.47 83.3 40.6 5.07 79.7 26.3 5.30 82.3 40.6 4.93 Dynamic+A 83.1 51.4 4.41 85.1 41.3 5.03 80.0 31.4 5.37 83.3 42.3 4.91 Dynamic+H 88.1 61.1 4.29 81.7 41.5 4.60 79.4 30.8 5.31 83.2 45.4 4.66 Deepseek-R1-32B Base 89.5 32.2 6.16 85.0 25.8 6.64 88.6 19.7 6.65 87.3 26.5 6.49 Static+A 92.3 45.8 5.14 92.5 35.8 5.39 91.5 20.8 5.99 92.2 35.6 5.45 Dynamic+A 92.6 51.4 5.06 93.3 38.0 5.33 91.4 21.9 5.93 92.6 38.6 5.38 Dynamic+H 93.5 51.2 4.82 92.4 38.1 5.23 87.8 33.2 5.46 91.7 41.2 5.15 Deepseek-V3 Base 93.7 66.3 5.26 93.8 39.7 6.00 94.7 31.1 6.22 94.0 46.3 5.81 Static+A 92.9 70.5 4.31 95.8 47.4 4.66 93.4 31.1 5.05 94.3 51.1 4.64 Dynamic+A 93.2 71.6 4.36 95.7 50.5 4.68 93.3 33.3 4.97 94.4 53.4 4.64 Dynamic+H 97.9 71.8 4.40 96.3 48.5 4.45 97.0 44.9 5.19 97.0 55.2 4.60 GPT-4o Base 92.0 62.7 5.07 91.0 35.2 5.67 94.5 27.8 6.26 92.1 42.3 5.61 Static+A 99.5 72.1 4.21 98.3 43.6 5.35 97.8 37.9 5.85 98.6 51.5 5.10 Dynamic+A 99.9 76.4 4.18 99.5 45.3 5.40 98.1 41.4 5.92 99.3 54.4 5.13 Dynamic+H 99.6 75.3 4.01 94.5 48.9 4.71 98.9 48.3 5.12 97.1 57.2 4.58 Qwen2.5-14B(SFT) Base 70.9 49.1 5.94 72.8 42.1 5.94 71.0 24.5 6.99 71.8 40.3 6.18 Static+A 84.6 61.4 4.50 78.1 38.6 4.69 77.8 35.6 5.65 80.1 45.3 4.85 Dynamic+A 85.8 64.9 4.58 78.4 39.9 4.75 78.1 39.0 5.59 80.7 47.7 4.89 Dynamic+H 82.7 64.6 4.59 81.4 48.9 4.67 78.5 37.4 5.74 81.2 51.3 4.89
Table 2: Impact of Naviagent Variants on ToolBench. Base retains only the core agent; StaticA augments with a static graph and Alpha-Beta pruning; DynamicA augments with a dynamic graph and Alpha-Beta pruning; DynamicH augments with a dynamic graph and heuristic pruning, which corresponds to our proposed NaviAgent. Metrics are reported as in Table 1. We also evaluate runtime, with detailed results reported in Appendix F.2.
@@ 5 Conclusion
We presented NaviAgent, a bilevel planning framework that separates high‑level decision making from low‑level execution over a tool world model, achieving robust gains on ToolBench and API‑Bank. It scales to thousands of tools with competitive efficiency and excels in complex, multi‑tool tasks and larger models. Remaining challenges include handling heterogeneous tool interfaces and dynamic conditions, which may be tackled via unified protocols and adaptive graph construction. Beyond tool reasoning, NaviAgent points to broader applications: by abstracting tools as agents, its evolving graph and decision space can naturally extend to multi‑agent collaboration. This perspective underscores both the challenges of building adaptive, robust systems and the opportunities for advancing toward more collaborative AI ecosystems.
@@ B.1 HGT Network
B.2 Graph Search Algorithm